OpenStudy (anonymous):

Systems of equations problem help! easy medal. Maya had $27. She spent all the money on buying 3 burgers for $x each and 2 sandwiches for $y each. If Maya had bought 2 burgers and 1 sandwich, she would have been left with $11. A student concluded that the price of each burger is $5 and the price of each sandwich is $6. Which statement best justifies whether the student's conclusion is correct or incorrec

OpenStudy (cj49):

3x+2y=27

OpenStudy (cj49):

2x+y=16

OpenStudy (cj49):

substitute the values as x=5 and y=6 and see if it satisfies the RHS

OpenStudy (anonymous):

A)The student's conclusion is correct because the solution to the system of equations 3x + 2y = 11 and 2x + y = 16 is (5, 6). B) The student's conclusion is incorrect because the solution to the system of equations 3x – 2y = 11 and 2x – y = 16 is (5, 6). c) The student's conclusion is correct because the solution to the system of equations 3x + 2y = 27 and 2x + y = 16 is (5, 6). D) The student's conclusion is incorrect because the solution to the system of equations 2x + 3y = 27 and x + 2y = 16 is (5, 6).

OpenStudy (cj49):

C

OpenStudy (anonymous):

thanks(:

OpenStudy (cj49):

np c: